Market research has revealed the following information about the market for chocolate




bars: The demand schedule can be represented by the equation QD = 1,600 – 300P, where




QD is the quantity demanded and P is the price. The supply schedule can be represented by




the equation QS = 1,400 + 700P, where QS is the quantity supplied. Calculate the equilibrium




price and quantity in the market for chocolate bars.

1
Expert's answer
2022-02-01T10:17:37-0500

Equilibrium Quantity and Price.

At equilibrium, the quantity demanded is equal to the quantity supplied.

"Q_d=1600-300P"

"Q_s=1400+700P"

"1600-300P=1400+700P"

collecting like terms together we have;

"1600-1400=700P+300P"

"200=1000P"

At equilibrium, the Price "\bold{P=0.20}"

with the equilibrium price, we substituted in the quantity demanded or supplied function to get the amount of quantity demanded at equilibrium.

"Q_d=1600-300(0.2)=1540"

"Q_s=1400+700(0.2)=1540"

Therefore the quantity at equilibrium is 1540.